OpenStudy (anonymous):

factor the expression of 9x2-4y2

OpenStudy (anonymous):

\[ x^2-49\]

OpenStudy (anonymous):

This is the difference of "2 perfect squares". -> x^2 is a square (obviously) and 49 is some particular integer squared. That part will be up to you to find it. And when you have 2 perfect squares: a^2 - b^2 that factors to (a - b)(a + b)

OpenStudy (anonymous):

@romanortiz65 do you know how to factor?

OpenStudy (anonymous):

isnt it (x-7)(x+7)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Good job! You got it!

OpenStudy (anonymous):

\[x^2 - 49 = (x+7)(x-7)\]

OpenStudy (anonymous):

All good now, @some_someone ?

OpenStudy (anonymous):

what about 9x2-4y2

OpenStudy (anonymous):

lol yeah i said he was right before :P

OpenStudy (anonymous):

I was just writing it again haha :P

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Anyway, 9x2-4y2 can be viewed as: (3x)^2 - (2y)^2 So, if you use the formula I wrote out for a = 3x and b = 2y you will get the answer.

OpenStudy (anonymous):

(x-3)(x-2)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(a - b)(a + b) = (3x - 2y)(3x + 2y)

OpenStudy (anonymous):

ok so for 4x2-4

OpenStudy (anonymous):

it would be

OpenStudy (anonymous):

(4x-4)(4x+4)

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

wtf damn i suck at this

OpenStudy (anonymous):

\[(4x-4)(4x+4) = 16x^2 + 16x - 16x - 16 \]

OpenStudy (anonymous):

(4x-2)(x+2)

OpenStudy (anonymous):

\[(4x-2)(x+2) = 4x^2 + 8x -2x - 4 = 4x^2 +6x - 4\]

OpenStudy (anonymous):

???????????????????

OpenStudy (anonymous):

Remember, you have to get these in the form of a^2 - b^2 first. 4x^2 - 4 = (2x)^2 - 2^2 = (a - b)(a + b) (2x - 2)(2x + 2) But your not done yet because you can further factor. 2(x - 1)2(x + 1) = (2)(2)(x - 1)(x + 1) = 4(x - 1)(x + 1) With the integer factors, you "bring them back together" to get the "4".

OpenStudy (anonymous):

so final answer would be 4(x-1)(x+1)

OpenStudy (anonymous):

Yes, that's correct. If you have a new one, just close this question and write it in a new post.